OpenStudy (anonymous):

I need help with this please. Solve using the addition and multiplication principles. 0.96y − 0.79 ≤ 0.21y + 0.46 = 0.96y + 0.21y – 0.79 < 0. 21y + 0.21y + 0.46 = 1.17y – 0.79 < 0.88y =0.38y < 0.88y Is this correct. If not how do i do the problem and what did I do wrong?

OpenStudy (anonymous):

You are close...you have the right idea.

OpenStudy (anonymous):

ok thanks. I dont no what step to do next. any tips.

OpenStudy (anonymous):

\[0.96y-0.79\le0.21y+0.46\]Move the 0.21y over to the left and the 0.79 over to the right\[0.96y-0.21y-0.79+0.79\le0.21y-0.21y+0.46+0.79\]\[0.75y\le1.15\]\[y\le(1.15)/(.75)=23/15=1.53333...\]

OpenStudy (anonymous):

That second line, i just subtracted 0.12y from both sides and added 0.79 to both sides to get all the y's together and all the non-y's together on opposite sides. Same method you do for equations.

OpenStudy (anonymous):

Sorry...subtracted 0.21y...

OpenStudy (anonymous):

ok cool. I got it know. Thanks for the help.

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

0.96y − 0.79 ≤ 0.21y + 0.46 = 0.96y - 0.21y – 0.79 < 0. 21y - 0.21y + 0.46 =0.04y < 0.46 = 0.04y/0.04 < 0.46/0.04 = y < 0.46/0.04 = y < 11.5 Check: 0.96y − 0.79 ≤ 0.21y + 0.46 Using 0 0.96() – 0.79 < 0.21(0) + 0.46 0.79 < 0.46 why did I get this answer?

OpenStudy (anonymous):

Because your problem is going from the 2nd line to the 3rd line. You subtracted 0.79 from the .96 as well as .21 0.79 is not a y-term and can't be combined with the 0.96y and the 0.21y. Instead, you need to add it to both sides. So, I'll start with your 2nd line and show you how it should be :) \[0.96y-0.21y-0.79 < 0.21y-0.21y+0.46\]\[0.75y-0.79+0.79<0.46+0.79\]\[0.75y<1.15\] Notice on my line 2, I combined the two terms that had a y in them. Then i moved the constant (0.79) from the left to the right. Does that help?